OpenStudy (mallorysipp234):

Given the parent functions f(x) = log10 x and g(x) = 5x − 2, what is f(x) • g(x)?

OpenStudy (mallorysipp234):

f(x) • g(x) = log10 x5x − 2 f(x) • g(x) = log10 (5x − 2)x f(x) • g(x) = 5x log10 x + 2 log10 x f(x) • g(x) = 2 log10 x − 5x log10 x

OpenStudy (***[isuru]***):

is it \[\log_{10} x \ or \ \log 10x \ \ ?\]

OpenStudy (***[isuru]***):

hint : for an example if u got \[f(x) = \log 7x\]\[g(x) = 2x + 2\] then \[f(x) . g(x) = \log7x( 2x + 2)\] so... for your problem it's \[f(x) = \log10x \]\[g(x) = 5x -2\] then according to the above example \[f(x) . g(x) = ?\]

OpenStudy (mallorysipp234):

then it would be f(x) * g(x) = log10x * 5x-2 ???

OpenStudy (***[isuru]***):

yep :)

OpenStudy (mallorysipp234):

alright so out of the choices i have then it would be B right? because parenthesis means multiplication correct?

OpenStudy (***[isuru]***):

yeah !

OpenStudy (mallorysipp234):

thank you! can you help me with a few more?

OpenStudy (mallorysipp234):

If f(x) = log 2 (x + 4) what is f ^ -1 (3) ??? the negative one is an exponent to the F.

OpenStudy (phi):

when you do function composition, you do f( g(x) ) which means you replace any x's in f() with g(x) if \( f(x)=\log_{10} x\) then \( f(g(x))=\log_{10} g(x)\) now replace g(x) with 5x − 2 \[ f(g(x))=\log_{10}(5x-2) \] notice that this is not x(5x-2) inside the log function

OpenStudy (phi):

unfortunately, I can not make sense of any of the answers.